# Improved Generation of Adversarial Examples Against Safety-aligned LLMs

### v1

@@ 3.3 Adapting Intermediate Level Attack for Gradient-base Attacks Against LLMs
Based on the observation, we adapt the strategy of ILA for use in gradient-based attacks adversarial prompt generation by maximizing the scalar projection of the last adversarial prompt token representation onto the directional guide. That is, we opt to maximize $L_{\mathrm{ILA}^{\dagger}}(x)=v_{r,n}^{T}h_{r,n}$. Another issue of the original ILA here is that it requires a preliminary attack, which is time-consuming to obtain in the setting of attacking LLMs. We make a compromise by utilizing the current adversarial prompt to obtain the directional guide, which leads to no increase in computational complexity. Note that we only modify the gradient computation step in each iteration. Therefore, the step of evaluating candidate adversarial suffixes can help reduce the adversarial loss, providing useful directional guidance in the initial iterations. We denote this adaptation of ILA as Language ILA (LILA) and evaluate its performance using GCG and AutoPrompt as baseline attacks against the Llama-2-7B-Chat model on AdvBench. As shown in Figure 7, LILA demonstrates a lower adversarial loss and a higher match rate compared to the baseline attacks. The experimental results suggest that our adaptation of ILA indeed improves discrete optimization in gradient-based adversarial prompt generation and offers insights to more effectively address similar discrete optimization problems within LLMs.
3.4 Combinations and Balance of Gradient Magnitude in Combinations
In the previous sections, we have confirmed the success of LSGM and LILA. To further improve the performance of discrete optimization using previous gradient-based attacks on LLMs, we then discuss possible approach to combining these strategies.
We first try to combine the LILA loss and the cross-entropy loss by introducing a scaling factor $\beta\geq 0$ which is used to balances the magnitude of the two loss terms, i.e., we minimize a combination loss $L_{\mathrm{CE}}-\beta L_{\mathrm{ILA}^{\dagger}}$. With this combination loss, the gradient w.r.t. $h_{r}$ can be written as $[\nabla_{h_{r,1}}L_{\mathrm{CE}}(x),...,\nabla_{h_{r,n}}L_{\mathrm{CE}}(x)-\beta v_{r},...,\nabla_{h_{r,n^{*}-1}}L_{\mathrm{CE}}(x)]$. It may result in excessively larger or smaller gradient flow through the last adversarial prompt token representation $h_{r,n}$ than other token representations, leading to overestimation or underestimation of the impact on adversarial loss resulting from the changes of the last adversarial prompt token representation. We alleviate the issue by re-normalizing the gradient w.r.t. the last adversarial prompt token representation, i.e., propagating $\alpha(\nabla_{h_{r,n}}L_{\mathrm{CE}}(x)-\beta v_{r})$ back to lower layers instead of using the original gradient, in which $\alpha=\|\nabla_{h_{r,n}}L_{\mathrm{CE}}(x)\|_{2}/\|\nabla_{h_{r,n}}L_{\mathrm{CE}}(x)-\beta v_{r}\|_{2}$. When we set $\beta\rightarrow\infty$, it can be approximately regarded as replacing the gradient w.r.t. $h_{r,n}$ with $-\alpha v$, and when we set $\beta=0$, it is equivalent to only using the cross-entropy loss.
(a) GCG
(b) AutoPrompt (AP)
Figure 8: Evaluating our combination method using (a) GCG and (b) AutoPrompt (AP) as baseline attacks. Best viewed in color.
We denote such a combination of the LILA loss and the cross-entropy loss as LILA. With LILA, it is natural to further combine it with LSGM and obtain LSGM-LILA. We evaluate using GCG and AutoPrompt as baseline attacks and report attack performance against Llama-2-7B-Chat on AdvBench, in the task of query-specific adversarial suffix generation. For simplicity, we can set $\beta\rightarrow\infty$ and adopt the aforementioned approximation for LILA. The results are shown in Table 8. It can be seen that by combining with the cross-entropy loss, LILA outperforms LILA, while LSGM-LILA achieves even higher match rates, demonstrating that the combination leads to obvious further improvement.
@@ 4.2 Experimental Results
Following the evaluations by GCG [52], we perform evaluations in the settings of query-specific adversarial suffix generation and universal adversarial suffix generation. For query-specific adversarial suffix generation, following [52], we use first 100 harmful behaviors in AdvBench. For universal adversarial suffix generation, we use the first 10 harmful queries in AdvBench to generate a universal adversarial suffix and test it on the rest 510 harmful queries in AdvBench. We use a Top-$k$ selection of 4 and a candidate set size of 20 for all methods at each iteration. Since Llama-2-Chat [41] models show great robust performance to gradient-based adversarial prompt generation [27], we mainly evaluate the methods on the model of Llama-2-7B-Chat [41] and the model of Llama-2-13B-Chat [41]. In addition, the model of Mistral-7B-Instruct-v0.2 [17] is also considered.
Table 1: Match rates, attack success rates, and time costs for generating query-specific adversarial suffixes on AdvBench are shown. The symbol indicates the use of the default setting for GCG, i.e., using a Top-$k$ of 256 and a candidate set size of 512. Time cost is derived by generating a single adversarial adversarial suffix on a single NVIDIA V100 32GB GPU.
Method Llama-2-7B-Chat Llama-2-13B-Chat Mistral-7B-Instruct Time cost
MR ASR MR ASR MR ASR
GCG* 60% 44% 58% 40% 95% 92% 85m
GCG 54% 38% 37% 33% 73% 74% 3m
GCG-LSGM 72% 62% 52% 43% 93% 88% 3m
GCG-LILA 62% 53% 38% 36% 75% 78% 3m
GCG-LILA 70% 59% 52% 48% 83% 80% 3m
GCG-LSGM-LILA 87% 68% 62% 52% 94% 93% 3m
@@
The comparison results in the setting of query-specific adversarial suffix generation are shown in Table 1. Experimental results demonstrate that both LSGM and LILA outperform the GCG attack on three safety-aligned LLMs. Our combination method further boosts both the match rates and the attack success rates, achieving the best performance. Specifically, GCG-LSGM-LILA achieves gains of $+30\%$, $+19\%$, and $+19\%$ when attacking the Llama-2-7B-Chat and Llama-2-13B-Chat, and Mistral-7B-Instruct, respectively. Moreover, since these methods reduce the gap between the input gradients and the effects of loss change results from token replacements, the size of candidate set at each iteration can be reduced for saving running time. We show the results of GCG with the default setting described in their paper, which evaluates 512 candidate adversarial suffixes at each iteration. It can be seen that our combination still shows outstanding performance, by only using $4\%$ time cost compared with the GCG with their initial setting (denoted as GCG in the table).
@@
Table 2: Attack success rates for generating universal adversarial suffixes on AdvBench. The average ASR (AASR), the worst ASR (WASR), and the best ASR (BASR) are obtained by performing each attack ten times.
Model GCG GCG-LSGM-ILA
@@
The results of the attacks in the setting of universal adversarial suffix generation are shown in Table 2. It can be observed that our combination not only achieves a remarkable improvement in the average ASR but also enhances both the worst and best ASRs obtained over 10 runs. Specifically, when attacking Llama-2-7B-Chat model, the GCG-LSGM-LILA achieves an average ASR of $60.32\%$, which gains a $+33.64\%$ improvement compared with the GCG attack. Moreover, for the worst and best ASR, GCG-LSGM-LILA achieves gains of $+34.82\%$ and $+31.06\%$, respectively.
@@ 5 Conclusions
In this paper, we present a new perspective on the discrete optimization problem in gradient-based adversarial prompt generation. That is, using gradient w.r.t. the input to reflect the change in loss that results from token replacement resembles using input gradient calculated on the substitute model to indicate the real effect of perturbing inputs on the prediction of a black-box victim model, which has been studied in transfer-based attacks against image classification models. By making some appropriate adaptations, we have appropriated the ideologies of two transfer-based methods, namely, SGM and ILA, into the gradient-based white-box attack against LLMs. Our analysis of the mechanisms behind their effective performance has provided new insights into solving discrete optimization problem within LLMs. Furthermore, we have developed an appropriate combination method to further enhance the discrete optimization in gradient-based adversarial prompt generation. Experimental results demonstrate that our combination method significantly improves the attack success rate.
@@ Appendix A Experimental Settings
Hyper-parameters. Unless further clarification is provided, we use a Top-$k$ selection of 4 and a candidate set size of 20 for all methods at each iteration. For SGM, we simply set $\gamma=0.5$. For LILA and LILA, we select the midpoint of the model layers. For instance, for Llama-2-13B-Chat [41], we select the 20-th layer, whereas for other models with 32 layers, we select the 16-th layer. We perform 500 iterations for all methods, with the number of adversarial tokens set to 20.
@@ Appendix B Ablation Study
(c) LILA
Figure 9: How the match rate and attack success rate change with (a) the choice of $\gamma$ for GCG-LSGM, (b) the choice of layer for GCG-LILA, and (c) the choice of layer for GCG-LILA. Best viewed in color.
We evaluate GCG-LSGM with varying the choice of $\gamma$ in Figure 9(a), GCG-LILA with varying the choices of the layer in Figure 9(b), and GCG-LILA with varying the choices of the layer in Figure 9(c), in the setting of query-specific adversarial suffix generation for attacking Llama-2-7B-Chat on AdvBench. It can be seen that for LSGM, a large range of the choice of $\gamma$ ($0.3\sim 0.9$) leads to improved performance. For our combination method, LILA demonstrates consistently more effective performance on all choices of the layer.

### v2

@@ 3.3 Adapting Intermediate Level Attack for Gradient-base Attacks Against LLMs
Another issue of the original ILA here is that it requires a preliminary attack, which is time-consuming to obtain in the setting of attacking LLMs. We make a compromise by utilizing the current adversarial prompt to obtain the directional guide, which leads to little increase in computational complexity. For instance, at $t$-th iteration, the directional guide is $v^{t}_{r,o}=h^{0}_{r,o}-h^{t}_{r,o}$. Note that we only modify the gradient computation step in each iteration. Therefore, the step of evaluating candidate adversarial suffixes can help reduce the adversarial loss, providing useful directional guidance in the initial iterations. We evaluate LILA using GCG and AutoPrompt as baseline attacks against the Llama-2-7B-Chat model on AdvBench. As shown in Figure 7, LILA demonstrates a lower adversarial loss and a higher match rate compared to the baseline attacks. The experimental results suggest that our adaptation of ILA indeed improves discrete optimization in gradient-based adversarial prompt generation and offers insights to more effectively address similar discrete optimization problems within LLMs.
@@ 4.2 Experimental Results
Following the evaluations by GCG [58], we perform evaluations in the settings of query-specific adversarial suffix generation and universal adversarial suffix generation. For query-specific adversarial suffix generation, following [58], we use first 100 harmful behaviors in AdvBench. For universal adversarial suffix generation, we use the first 10 harmful queries in AdvBench to generate a universal adversarial suffix and test it on the rest 510 harmful queries in AdvBench. We also test our method for universal adversarial suffix generation on the standard behaviors set from HarmBench [32], following the setting suggested in HarmBench. We use a Top-$k$ selection of 4 and a candidate set size of 20 for Llama and Mistral models, and a Top-$k$ selection of 64 and a candidate set size of 160 for Phi3-Mini-4K-Instruct. Since Llama-2-Chat [46] models show great robust performance to gradient-based adversarial prompt generation [32], we mainly evaluate the methods on the model of Llama-2-7B-Chat [46] and the model of Llama-2-13B-Chat [46]. In addition, Mistral-7B-Instruct-v0.2 [19] and Phi3-Mini-4K-Instruct [1] are also considered.
@@
The comparison results in the setting of query-specific adversarial suffix generation are shown in Table 1. Experimental results demonstrate that both LSGM and LILA outperform the GCG attack on three safety-aligned LLMs. Our combination method further boosts both the match rates and the attack success rates, achieving the best performance. Specifically, GCG-LSGM-LILA achieves gains of $+30\%$, $+19\%$, $+19\%$, and $+21\%$ when attacking the Llama-2-7B-Chat and Llama-2-13B-Chat, Mistral-7B-Instruct, and Phi3-Mini-4K-Instruct, respectively. Moreover, since these methods reduce the gap between the input gradients and the effects of loss change results from token replacements, the size of candidate set at each iteration can be reduced for saving running time. We show the results of GCG with the default setting described in their paper, which evaluates 512 candidate adversarial suffixes at each iteration. It can be seen that our combination still shows outstanding performance, by only using $4\%$ time cost compared with the GCG with their initial setting (denoted as GCG in the table).
Table 1: Match rates, attack success rates, and time costs for generating query-specific adversarial suffixes on AdvBench are shown. The symbol indicates the use of the default setting for GCG, i.e., using a Top-$k$ of 256 and a candidate set size of 512. Time cost is derived by generating a single adversarial suffix on a single NVIDIA V100 32GB GPU.
Method Llama-2-7B-Chat Llama-2-13B-Chat Mistral-7B-Instruct Phi3-Mini-4K-Instruct
MR ASR Time MR ASR Time MR ASR Time MR ASR Time
GCG* 60% 44% 85m 58% 40% 170m 95% 92% 85m 70% 61% 50m
GCG 54% 38% 3m 37% 33% 6m 73% 74% 3m 60% 59% 17m
GCG-LSGM 72% 62% 3m 52% 43% 6m 93% 88% 3m 75% 64% 17m
GCG-LILA 70% 59% 3m 52% 48% 6m 83% 80% 3m 65% 61% 17m
GCG-LSGM-LILA 87% 68% 3m 62% 52% 6m 94% 93% 3m 81% 68% 17m
The results of the attacks in the setting of universal adversarial suffix generation are shown in Table 2. It can be observed that our combination not only achieves a remarkable improvement in the average ASR but also enhances both the worst and best ASRs obtained over 10 runs. Specifically, when attacking Llama-2-7B-Chat model on AdvBench, the GCG-LSGM-LILA achieves an average ASR of $60.32\%$, which gains a $+33.64\%$ improvement compared with the GCG attack. Moreover, for the worst and best ASR, GCG-LSGM-LILA achieves gains of $+34.82\%$ and $+31.06\%$, respectively.
@@
Table 2: Attack success rates for generating universal adversarial suffixes on AdvBench and HarmBench. The average ASR (AASR), the worst ASR (WASR), and the best ASR (BASR) are obtained by performing each attack ten times.
Dataset Model GCG GCG-LSGM-ILA
@@
Phi3-Mini-4K-Instruct 35.84% 20.39% 46.27% 44.80% 31.18% 61.18%
HarmBench Llama-2-7B-Chat 56.90% 33.00% 66.50% 69.35% 57.50% 87.00%
Llama-2-13B-Chat 37.40% 13.50% 64.50% 53.55% 22.00% 81.50%
Mistral-7B-Instruct 75.00% 37.50% 90.50% 81.00% 66.50% 93.00%
Phi3-Mini-4K-Instruct 49.90% 29.00% 65.50% 63.80% 48.50% 80.50%
We also test the transfer attack performance of our method. We use the universal suffixes generated by performing GCG and GCG-LSGM-LILA against Llama-2-7B-Chat to attack GPT-3.5-Turbo on 100 harmful queries in AdvBench. The test queries are distinct from the queries that are used to generate universal suffixes. The results are shown in Table 3. It can be observed that our GCG-LSGM-LILA achieves remarkable improvements in the average, worst, and best ASRs obtained over 10 runs.
Table 3: The performance of transfer attack against GPT-3.5-Turbo on AdvBench. The average ASR (AASR), the worst ASR (WASR), and the best ASR (BASR) are obtained by performing each attack ten times.
Method AASR WASR BASR
GCG 38.30% 24% 48%
GCG-LSGM-LILA 45.20% 35% 81%
@@ 5 Conclusions
In this paper, we present a new perspective on the discrete optimization problem in gradient-based adversarial prompt generation. That is, using gradient w.r.t. the input to reflect the change in loss that results from token replacement resembles using input gradient calculated on the substitute model to indicate the real effect of perturbing inputs on the prediction of a black-box victim model, which has been studied in transfer-based attacks against image classification models. By making some appropriate adaptations, we have appropriated the ideologies of two transfer-based methods, namely, SGM and ILA, into the gradient-based white-box attack against LLMs. Our analysis of the mechanisms behind their effective performance has provided new insights into solving discrete optimization problem within LLMs. Furthermore, the combination of these methods can further enhance the discrete optimization in gradient-based adversarial prompt generation. Experimental results demonstrate that our methods significantly improves the attack success rate for both white-box and transfer attacks.
@@ Appendix A Experimental Settings
Hyper-parameters. For SGM, we simply set $\gamma=0.5$. For the selection of intermediate representation to perform LILA, we choose the first token from the target phrase since it empirically performs better than selecting other tokens or all tokens except the adversarial suffix and the last input token. For the intermediate layer, we choose the midpoint of the model layers. For instance, for Llama-2-13B-Chat [46], we select the 20-th layer, whereas for other models with 32 layers, we select the 16-th layer. We perform 500 iterations for all methods, with the number of adversarial tokens set to 20.
@@ Appendix B Ablation Study
Figure 8: How the match rate and attack success rate change with (a) the choice of $\gamma$ for GCG-LSGM, (b) the choice of layer for GCG-LILA. Best viewed in color.
We evaluate GCG-LSGM with varying the choice of $\gamma$ in Figure 8(a), and GCG-LILA with varying the choices of the layer in Figure 8(b), in the setting of query-specific adversarial suffix generation for attacking Llama-2-7B-Chat on AdvBench. It can be seen that for LSGM, a large range of the choice of $\gamma$ ($0.3\sim 0.9$) leads to improved performance. For LILA, it demonstrates consistently more effective performance on all choices of the layer.